Gerry is twice Julia's age.Ten years ago,the square of Julia's age was 14 more than Gerry's age.What is the present age of Gerry?
Answer by stanbon(75887)

Gerry is twice Julia's age.Ten years ago,the square of Julia's age was 14 more than Gerry's age.What is the present age of Gerry?
-------
Equations:
G = 2J
(J-10)^2 = (G-10)+14
----
Substitute and solve for "J":
(J-10)^2 = (2J-10) + 14
J^2 - 20J + 100 = 2J + 4
J^2 - 22 J + 96 = 0
J = 16
---
Solve for G:
G = 2J = 32 (Gerry's age now)
------------
Cheers,
Stan H.